A tennis ball connected to a string is spun around in a vertical, circular path at...

tennisball

A tennis ball connected to a string is spun around in a vertical, circular path at a uniform speed. The ball has a mass m = 0.153 kg and moves at v = 5.08 m/s. The circular path has a radius of R = 0.99 m

1)What is the magnitude of the tension in the string when the ball is at the bottom of the circle?

2)What is the magnitude of the tension in the string when the ball is at the side of the circle?

3)What is the magnitude of the tension in the string when the ball is at the top of the circle?

4)What is the minimum velocity so the string will not go slack as the ball moves around the circle?
Answer #1

mass of ball = 0.153kg

weight = 0.153*9.8 = 1.4994N

velocity = 5.08m/s

R = 0.99m

centrifugal force = F = mv2/R = 3.9883N

1)

at the bottom,

T = F + weight = 5.4877N

2)

At the side = T = sqrt(F2 + weight2) = 4.2608N

3)

at the top,

T = F - weight = 2.4889

4)

minim speed = vm = ?

mv2/R = mg

v = sqrt(gR) = 3.1148m/sec